# Multi-Draft Speculative Sampling: Canonical Architectures and Theoretical Limits

### v1

@@ 4 Faster Importance Weighted Speculative Sampling
In practice the distribution of the target and draft model is often concentrated over a small number of tokens. It has also been observed that sampling from a high probability set (such as the top-$k$ highest probability tokens) or the top-$p$ set (the set of high probability tokens with aggregate probability exceeding a threshold) leads to more coherent outputs (Meister et al., 2023). After such top-$p$ sampling the effective alphabet size, i.e., the number of tokens with non-zero probability is generally small. We report the histogram of effective alphabet size for the OPT model in Appendix F. This motivates us to develop some approaches for speeding up our proposed solution by reducing the number of variables required in optimization.
We focus on the case of $K=2$ drafts and explain how to extend the approach to the general case. Let $X_{1}$ and $X_{2}$ denote the input tokens and $Y$ denote the selected token. When $X_{1}=X_{2}=i$ we have that $\beta_{i}(i,i)=1$. Furthermore as discussed previously, due to symmetry we have $\beta_{y}(i,j)=\beta_{y}(j,i)$. It is more convenient to introduce a new set of variables $w_{i,j}$ which are defined as:
@@
i.e., $w_{i,j}$ denotes the probability that the output token is $i$ given that we see the pair $\{i,j\}$ at the input in any order. For any given choice of $w_{i,j}$ the distribution of $Y$ can be computed as follows:
$p_{I}(k)=\Pr(Y=k)=p_{k}^{2}+\sum_{i=1,i\neq k}^{n}2p_{i}p_{k}w_{i,k}$ (11)
where we use ${\mathbf{p}}=(p_{1},\ldots,p_{n})$ as the probability vector for the draft model and ${\mathbf{q}}=(q_{1},\ldots,q_{n})$ as the probability vector for the target model. Our aim is to maximize the following objective $\sum_{i=1}^{n}\min\left(p_{I}(i),q_{i}\right)$ over the variables $w_{i,j}$ that satisfy $0\leq w_{i,j}\leq 1$ and $w_{i,j}+w_{j,i}=1$. This optimization can be cast as a linear programming problem with $O(n^{2})$ variables which may be slow in practice. With this formulation as a starting point we discuss two approaches to reduce the number of variables needed to optimize.
Our first observation is that the linear programming solution will not be sensitive to most choices of $w_{i,j}$ when the target and draft distributions are concentrated over a few values. As a result one can heuristically set most of the variables. We refer to this method as the truncated LP scheme and present the method in Algorithm 1. Assume that the vocabulary $\Omega=\{1,2,\ldots,n\}$ has the tokens sorted in decreasing order i.e., $q_{1}-p_{1}^{2}\geq q_{2}-p_{2}^{2}\ldots\geq q_{n}-p_{n}^{2}$. We partition $\Omega$ into two sets $\Omega_{1}=\{1,2,\ldots,s\}$ and $\Omega_{2}=\{s+1,\ldots,n\}$, where $s$ is a free parameter to select. We fix a subset of weights as follows:
@@
while we leave the weights $w_{i,j}$ for $i<j$ and $i,j\in\Omega_{1}$ as free parameters. The intuition behind the choice of weights in (12) is that in these cases we prefer token $i$ over token $j$ to increase $p_{I}(i)$ further so as to decrease the difference between $q_{i}$ and $p_{I}(i)$. Note that:
@@
The objective to maximize reduces to $\sum_{k=1}^{s}\min(p_{I}(k),q_{k})$ over the variables $w_{i,j}$. Thus the number of variables is reduced to $O(s^{2})$. We further show in Appendix E that if $P^{\star}(\mathrm{acc})$ is the optimal acceptance probability associated by applying the linear program over all $O(n^{2})$ weight variables and $\tilde{P}(\mathrm{acc})$ is the acceptance probability for the truncated program then:
@@
Algorithm 1 Truncated LP
2:  Output: Selected token $Y_{I}$, output distribution $p_{I}(\cdot)$.
4:  Set $\Omega_{1}=\{1,\ldots,s\}$ and $\Omega_{2}=\{s+1,\ldots,n\}$.
  • Constraints: $w_{i,j}\geq 0$, $w_{i,j}+w_{j,i}=1$
7:  Compute $p_{I}(\cdot)$ according to (13).
8:  if $X_{1}=X_{2}$ then
9:     Set $Y_{I}=X_{1}$.
10:  else
11:     Let $\{X_{1},X_{2}\}=\{i,j\}$ and $i<j$.
12:     Set $Y_{I}=i$ with probability $w_{i,j}$, and $Y_{I}=j$ otherwise.
13:  end if
Thus if $\Omega_{2}$ is selected so that the penalty term is small then the decrease in the acceptance probability can be kept small. In the experiments we observed that for well-trained target models the drop in accuracy is negligible even for small values of $s$. Thus by appropriately truncating the number of variables to optimize in the linear program we expect to have a faster implementation.
Our second idea is to truncate the alphabet when performing importance sampling. In particular let $\Omega_{0}\subseteq\Omega$ be a high probability subset of $\Omega$ under the target distribution. We re-normalize the target distribution to $\tilde{q}$ so that it is supported entirely over $\Omega_{0}$ and use this distribution in our proposed scheme to generate an output $Y\sim\tilde{q}(\cdot)$. Since we want the output to follow the distribution $q(\cdot)$, we perform the following post-processing. We accept $Y$ with a probability $p_{a}=\sum_{x\in\Omega_{0}}q(x)$ and reject with probability $p_{r}=1-p_{a}$. If rejected, we output a token from $\Omega\setminus\Omega_{0}$ such that any $x\in\Omega\setminus\Omega_{0}$ is selected with probability proportional to $q(x)$. This ensures that the output token follows the original distribution $q(\cdot)$. We refer to this scheme as the truncated alphabet scheme.
Remark 3
Although our discussion has focused on the case when both samples are drawn from the same draft distribution $p(\cdot)$, our approach can also be extended to the case when they are sampled from a different distribution, as outlined in the supplementary material in Section G.
@@ 5 Experimental Results
Setup. We conduct experiments using an instance of A100 GPU with 80GB memory. We use the OPT model (Zhang et al., 2022), where the draft model has $125$M parameters and the target model has $13$B parameters. For evaluation purposes, we consider the datasets associated with the XSum (Narayan et al., 2018), Databricks-Dolly-15k (Conover, Mike and Hayes and others, 2023) and the WMT18 (Bojar et al., 2018) tasks. All experiments were done over $4$ arbitrarily chosen seeds and the performance was averaged over these.
Parameters. In applying our Importance Sampling (IS) scheme to handle $K>2$ drafts, we utilize the successive selection scheme discussed in Remark 4. Throughout our experiments, unless specified otherwise, we employ both truncated LP (with $s=5$ as the truncation parameter) and truncated alphabet (to a size of $40$ tokens) as detailed in Section 4. Additionally, we perform top-$p$ sampling with $p=0.95$ and generate $5$ tokens per call of the draft model across all schemes. Further experimental results, including empirical evaluations of acceptance probabilities and results with top-$k$ sampling, are presented in Appendix Section H.
5.1 Identical Draft Models
In our first set of experiments, we consider the case where identical draft models are used to generate candidate tokens sequences. We compare the performance of our method with SpecTr (Sun et al., 2024b) and SpecInfer (Miao et al., 2024), as well as single-draft speculative sampling (Leviathan et al., 2023; Chen et al., 2023), where we use the same temperature for the draft model.
In Fig. 3, we set the temperature of the target model to $1.0$, while we vary the temperature of the $K=2$ draft models between the range of $1.2$ to $2.4$, and we report the performance achieved by the different schemes across the three tasks discussed earlier. The top plots report the block efficiency, which is the average number of tokens that are accepted per use of the draft model (Leviathan et al., 2023). The bottom plots show the percentage improvement in the token rate with respect to the baseline single draft scheme.
We observe that the IS scheme consistently outperforms the both SpecTr and SpecInfer across all three tasks. In fact when the temperature of the second draft is increased, the improvement in the block efficiency of both SpecTr and SpecInfer is rather negligible compared to the single draft baseline, while the token rate are in fact lower than the single draft baseline. On the other hand our proposed IS scheme is able to achieve consistent improvements in all the three tasks. In the experiments involving the XSum and WMT tasks we also measure the ROUGE-L (Lin, 2004) and BLEU (Papineni et al., 2002) scores respectively, which are reported in the Appendix H.3.
Scheme $K=2$ $K=3$ $K=4$ $K=5$ $K=6$
IS 2.13 $\pm$ 0.05 2.22 $\pm$ 0.05 2.26 $\pm$ 0.05 2.27 $\pm$ 0.05 2.28 $\pm$ 0.06
5.2 Non-Identical Draft Models
We next consider the case of $K=2$ draft models, which share the same weights but use different temperatures for token generation. When the candidate tokens are generated from different distributions, to the best of our knowledge, the only multi-draft sampling scheme that has been considered in prior works is the SpecInfer scheme. As a result we are able to provide a direct comparison with this scheme. We set the temperature of the target model to $1.0$, and one that of the draft models to $1.2$ while we vary the temperature of the other draft model between the range of $1.0$ to $2.4$.
@@ 5.2 Non-Identical Draft Models
Figure 4: Performance comparison of different multi-draft schemes. The temperature of the first draft model is set to $1.2$, while we vary the temperature of the other draft.
In Fig. 4 we report the performance achieved by the different schemes across three different datasets. We observe that the IS scheme consistently outperforms SpecInfer across all three tasks. In fact when the temperature of the second draft is increased, the improvement in the block efficiency for SpecInfer is rather negligible compared to the single draft baseline. On the other hand our proposed IS scheme is able to achieve consistent improvements.
As with the case of identical drafts, we again observe that when the temperature of the second draft increases, the gains are negative for the SpecInfer scheme as the computational time for the second draft does not translate into a sufficient improvement in the block efficiency. On the other hand our proposed importance sampling scheme shows a consistent improvement the token-rate over the baseline schemes due to improved block efficiency, despite the additional compute from the second draft. As before, we report the ROUGE-L and BLEU scores in the Appendix H.3.
@@
Block Efficiency Token Rate
(improvement to SD)
10 1.98 $\pm$ 0.03 -0.57 $\pm$ 3.38%
Alphabet Truncation ($|\Omega_{0}|$) 20 2.00 $\pm$ 0.04 1.00 $\pm$ 3.08%
@@
In Table 2 we study the effect of choosing different parameters for vocabulary truncation and LP truncation discussed in Section 4. We use the same parameters as in the previous experiment but fix the temperature of the second draft to $2.0$. As we increase the size of the vocabulary $\Omega_{0}$ from $10$ to $40$ we see improvements in the block efficiency as well as the token rates. Beyond $40$, it appears that the gains in the block efficiency saturate and the token rate decreases. We also find that the performance is not too sensitive to the choice of the parameter $s$ in the LP program and choosing $s=5$ yields the best token rate.
@@ 6 Conclusion
We revisit the problem of maximizing the acceptance probability for token-level multi-draft selection (Sun et al., 2024b). We first present an architectural result that demonstrates connection to importance sampling. For the case of $K=2$ drafts we establish a closed for solution for the acceptance probability and also present a necessary and sufficient condition for the acceptance probability to equal one. Building on our theoretical analysis, we present a new class of token selection schemes based on importance sampling. We present experimental results using OPT model and three tasks: Dolly, XSum and WMT to compare the performance of these schemes over baselines. In particular we noted that our proposed schemes lead to a noticeable improvement in the block efficiency and token rates when the draft sequences are sampled from non-identical distributions.
Appendix A Proof of Theorem 1

### v2

@@ 4 Faster Importance Weighted Speculative Sampling
We discuss fast approaches for computing $\beta_{y}(\cdot)$ in our canonical decomposition in Fig. 1. Ideally we wish to select $\beta_{y}(\cdot)$ that maximize the acceptance probability in (2), but this involves a computationally expensive linear program. Fortunately we can develop procedures for faster computation that still achieve high acceptance probability. In practice the distribution of the target and draft model is often concentrated over a small number of tokens. It has also been observed that sampling from a high probability set such as the top-$p$ set (the set of high probability tokens with aggregate probability exceeding a threshold) leads to more coherent outputs (Meister et al., 2023). After such sampling the effective alphabet size, i.e., the number of tokens with non-zero probability, is generally small. We report some measurements of the effective alphabet size for the OPT model in Appendix H. This motivates us to develop some approaches for speeding up our proposed solution by reducing the number of variables required in optimization. Throughout this section we consider the case when the drafts have identical distribution and discuss the case of non-identical draft distributions in Appendix I. We focus on the case of $K=2$ drafts and then discuss how to tackle the general case.
Linear Program for $K=2$ drafts: We first consider the case of $K=2$ drafts and revisit the linear program that needs to be solved to compute the optimal acceptance probability in (2). We will then explain how to reduce the variables in the linear program for faster computation. Let $X_{1}$ and $X_{2}$ denote the input tokens and $Y$ denote the selected token. Note that when $X_{1}=X_{2}=i$, we have that $\beta_{i}(i,i)=1$. Furthermore when the draft models have identical distributions, from symmetry, we have $\beta_{y}(i,j)=\beta_{y}(j,i)$. It is more convenient to introduce a new set of variables $w_{i,j}$ which are defined as:
@@
i.e., $w_{i,j}$ denotes the probability that the output token is $i$ given that we see the pair $\{i,j\}$ at the input in any order. We next discuss how to formulate a linear program involving the variables $w_{i,j}$ to maximize the acceptance probability in (2). For convenience we let ${\mathbf{p}}=(p_{1},\ldots,p_{n})$ as the probability vector for the draft model and ${\mathbf{q}}=(q_{1},\ldots,q_{n})$ as the probability vector for the target model. The acceptance probability i.e., the objective in (2), is given by $\sum_{i=1}^{n}\min\left(p_{I}(i),q_{i}\right)$, where
$p_{I}(k)=p_{k}^{2}+\sum_{i=1,i\neq k}^{n}2p_{i}p_{k}w_{k,i}$ (9)
denotes probability that the selected token is $Y=k$. Furthermore for our definition of $w_{i,j}$ in (8), the associated constraints are: $w_{i,j}+w_{j,i}=1$ and $0\leq w_{i,j}\leq 1$. This optimization can be cast as a linear programming problem with $O(n^{2})$ variables which may be slow in practice. We next discuss techniques to reduce the number of variables in optimization without a significant loss in the acceptance probability.
Faster Approximations to Linear Program: In order to propose a faster approximation, we note that when the draft and target distribution are concentrated over a few tokens, the linear programming solution will not be sensitive to most choices of $w_{i,j}$. As a result one can heuristically set most of the variables. We refer to this method as the truncated LP scheme and present the pseudocode in Algorithm 2 in Appendix G. Assume that the vocabulary $\Omega=\{1,2,\ldots,n\}$ has the tokens sorted in decreasing order i.e., $q_{1}-p_{1}^{2}\geq q_{2}-p_{2}^{2}\ldots\geq q_{n}-p_{n}^{2}$. We partition $\Omega$ into two sets $\Omega_{1}=\{1,2,\ldots,s\}$ and $\Omega_{2}=\{s+1,\ldots,n\}$, where $s$ is a design parameter to select. We fix a subset of weights as follows:
@@
while we leave the weights $w_{i,j}$ for $i<j$ and $i,j\in\Omega_{1}$ as free parameters. The intuition behind the choice of weights in (10) is that in these cases we prefer token $i$ over token $j$ to increase $p_{I}(i)$ further, which is in turn can decrease the difference between $q_{i}$ and $p_{I}(i)$. Note that (9) reduces to:
@@
Our objective is to maximize reduces to $\sum_{k=1}^{s}\min(p_{I}(k),q_{k})$ over the variables $w_{i,j}$. Thus the number of variables is reduced to $O(s^{2})$. We further show in Appendix F that if $P^{\star}(\mathrm{acc})$ is the optimal acceptance probability associated by applying the linear program over all $O(n^{2})$ weight variables and $\tilde{P}(\mathrm{acc})$ is the acceptance probability for the truncated program then:
@@
Thus if $\Omega_{2}$ is selected so that the penalty term is small then the decrease in the acceptance probability can be kept small.
In the experiments we observed that for well-trained target models the drop in accuracy is negligible even for small values of $s$. Thus by appropriately truncating the number of variables to optimize in the linear program we expect to have a faster implementation. We discuss the computational complexity of the proposed method in Appendix G.
Our second proposal is to truncate the alphabet when performing importance sampling. In particular let $\Omega_{0}\subseteq\Omega$ be a high probability subset of $\Omega$ under the target distribution. We re-normalize the target distribution to $\tilde{q}$ so that it is supported entirely over $\Omega_{0}$ and use this distribution in our proposed scheme to generate an output $Y\sim\tilde{q}(\cdot)$. Since we want the output to follow the distribution $q(\cdot)$, we perform the following post-processing. We accept $Y$ with a probability $p_{a}=\sum_{x\in\Omega_{0}}q(x)$ and reject with probability $p_{r}=1-p_{a}$. If rejected, we output a token from $\Omega\setminus\Omega_{0}$ such that any $x\in\Omega\setminus\Omega_{0}$ is selected with probability proportional to $q(x)$. This ensures that the output token follows the original distribution $q(\cdot)$. We refer to this scheme as the truncated alphabet scheme.
@@ 5 Experimental Results
Setup. We conduct experiments using an instance of A100 GPU with 80GB memory. We use the OPT models (Zhang et al., 2022), where the draft model has $125$ million parameters and the target model has $13$B parameters. For evaluation purposes we consider the datasets associated with the XSum (Narayan et al., 2018), Databricks-Dolly-15k (Conover et al., 2023) and the WMT18 (Bojar et al., 2018) tasks. All experiments were done over $4$ arbitrarily chosen seeds and the performance was averaged over these.
5.1 Experiments with top-$p$ sampling
@@
Figure 3: Performance comparison of different multi-draft schemes. The temperature of the first draft models is set to $1.2$, while we vary the temperature of the other draft.
In this section we report experiments on LLM tasks when top-$p$ sampling is used with $p=0.95$. In our first set of experiments in Fig. 3, we consider the case of $K=2$ draft models, which share the same weights but use different temperatures for token generation. We set the temperature of the target model to $1.0$, and one that of the draft models to $1.2$ while we vary the temperature of the other draft model between the range of $1.0$ to $2.4$. In all our experiments we generate $5$ tokens per call of the draft model.
Our baseline scheme is the single-draft speculative sampling scheme (Leviathan et al., 2023, Chen et al., 2023) when we only use a single draft with sampling temperature of $1.2$. The other two schemes are the SpecInfer (Miao et al., 2024) and our proposed importance sampling (IS) scheme. In the IS scheme we employ both truncated LP (with $s=5$ as the truncation parameter) and truncated alphabet (to a size of $40$ tokens) as discussed in section 4. In Fig. 3 we report the performance achieved by the different schemes across the three tasks. The top plots report the block efficiency, which is the average number of tokens that are accepted per use of the draft model (Leviathan et al., 2023). The block efficiency achieved by the single-draft scheme is shown by the horizontal dotted red line in each plot, while the block efficiency of the IS scheme and SpecInfer are shown using the dark blue and light blue bars. We observe that the IS scheme consistently outperforms SpecInfer across all three tasks. In fact when the temperature of the second draft is increased, the improvement in the block efficiency for SpecInfer is rather negligible compared to the single draft baseline. On the other hand our proposed IS scheme is able to achieve consistent improvements. The bottom plots show the percentage improvement in the token rate with respect to the baseline single draft scheme. We observe that when the temperature of the second draft increases, the gains are negative for the SpecInfer scheme as the computational time for the second draft does not translate into a sufficient improvement in the block efficiency. On the other hand our proposed importance sampling scheme shows a consistent improvement the token-rate over the baseline schemes due to improved block efficiency, despite the additional compute from the second draft. In the experiments involving the XSUM and WMT tasks we also measure the ROUGE-L (Lin, 2004) and BLEU (Papineni et al., 2002) scores respectively, which are reported in the Appendix J.3. In Appendix J.1 we provide a similar experiment when both drafts use identical temperatures. We are able to also include comparisons with the SpecTr scheme (Sun et al., 2024a) in that experiment. Likewise in Appendix J.2 we present a similar experiment involving $K=3$ drafts.
In Table 1 we study the effect of choosing different parameters for vocabulary truncation and LP truncation discussed in Section 4. We use the same parameters as in the previous experiment but fix the temperature of the second draft to 2.0. As we increase the size of the vocabulary $\Omega_{0}$ from $10$ to $40$ we see improvements in the block efficiency as well as the token rates. Beyond $40$, it appears that the gains in the block efficiency saturate and the token rate decreases. We also find that the block efficiency is not too sensitive to the choice of the parameter $s$ in the LP program and choosing $s=5$ yields the best token rate.
@@
Block Efficiency Token Rate % improvement to SD
Vocab. Truncation ($|\Omega_{0}|$) 10 1.98 $\pm$ 0.03 -0.57 $\pm$ 3.38%
20 2.00 $\pm$ 0.04 1.00 $\pm$ 3.08%
@@
In Figure 4, we fix the temperature of the two draft models to $1.0$, and vary the temperature the target model in the range of $0.2$ to $1.0$. The rest of the parameters remain the same. We again report the block efficiency and the improvement in the token rate over the single draft scheme with the same draft temperature over the Dolly task. As the two drafts have identical temperature we are also able to include comparisons with the SpecTr scheme. We again observe that our proposed method generally attains the best performance.
Figure 4: Performance comparison of different schemes on the Dolly task, while we vary the temperature of the target model and keeping the temperature of the two drafts to $1.0$.
5.2 Experiments with top-$k$ sampling
We next consider experiments when top-$k$ sampling is applied to LLM tasks. We again use the same OPT models as in the previous subsection for both draft and target models. As before, we also sample a total of $5$ tokens during each call to the draft model. In the Dolly task we use a sampling temperature of $1.0$ for both the target and draft models. In the XSUM task we use a temperature of $0.9$ for the target and a temperature of $0.3$ for the draft model.
In the first experiment, reported in Table 2, we compare the toke-level acceptance probability across different methods. We first generate a sequence of tokens via auto-regressive decoding using the target model. At each step we compute the logits generated by the draft and the target model, which we then use to compute the acceptance probability for different schemes. We average the acceptance probability over $100$ instances. In the importance sampling (IS) scheme, we use a truncated linear program with a cut-off threshold $s=5$. We also consider the theoretically optimal acceptance probability for the case of $K=2$ drafts and also compute the natural extension of this expression for $K=4,8$ discussed in Remark 2 (reported using the gray font). We use top-$k$ sampling with $k=5$ for both models in this experiment. We report the acceptance probabilities for the case of $K=2,4,8$ drafts. In applying the IS scheme to handle $K>2$ drafts, we use the successive selection scheme discussed in Remark 3. Note that the acceptance probability increases as we increase the number of drafts in each case.
Table 2: Comparison of average acceptance probability across different tasks.
Scheme XSUM Dolly
$K=2$ $K=4$ $K=8$ $K=2$ $K=4$ $K=8$
Optimal 0.5009 0.5226 0.5419 0.6384 0.6731 0.6962
IS 0.4933 0.5145 0.5333 0.6348 0.6691 0.6919
SpecTr 0.4889 0.5083 0.5263 0.6246 0.6560 0.6800
SpecInfer 0.4875 0.5058 0.5227 0.6202 0.6489 0.6722
In Table 3 we compare the block efficiencies for different methods using $K=2$ and $K=3$ drafts. We apply top-$k$ sampling with $k=10$ and $k=5$ and use a temperature of $1.0$ for both models. In our IS scheme, we truncate the variables in the LP to $s=7$. While our proposed scheme still achieves a higher block efficiency over prior works, the gains are relatively small. This can be explained by noting that in this setting the acceptance probability of the baseline schemes is already close to the theoretical limit, as noted in Table 2.
Sampling Scheme $K=2$ drafts $K=3$ drafts
Block Efficiency Loss Block Efficiency Loss
top-$k$ ($k=10$) IS $2.48\pm 0.01$ $2.59\pm 0.02$
SpecTr $2.43\pm 0.01$ $98\%$ $2.55\pm 0.01$ $98\%$
SpecInfer $2.38\pm 0.02$ $96\%$ $2.49\pm 0.02$ $96\%$
top-$k$ ($k=5$) IS $2.52\pm 0.02$ $2.63\pm 0.03$
SpecTr $2.48\pm 0.02$ $98\%$ $2.56\pm 0.03$ $97\%$
SpecInfer $2.47\pm 0.01$ $98\%$ $2.55\pm 0.04$ $97\%$
@@ 6 Conclusion
We revisit the problem of maximizing the acceptance probability for token-level multi-draft selection (Sun et al., 2024b). We first present a decomposition result that demonstrates connection to importance sampling. For the case of $K=2$ (identical) drafts we establish a closed for solution for the acceptance probability and also present a necessary and sufficient condition for the acceptance probability to equal one. We propose a practical token selection scheme that mimics the theoretical decomposition principle and provides a way to tradeoff the computational speed with the acceptance probability. We present experimental results using OPT model and three tasks: Dolly, XSUM and WMT and demonstrate improvements in block efficiency and token rates over baselines schemes.
Acknowledgments
We thank Mingu Lee, Wonseok Jeon, and Babak Ehteshami Bejnordi for their help with this work.
Appendix A Speculative Decoding – Background
We briefly review the proposal made in Leviathan et al. (2023), Chen et al. (2023). Let $\Omega$ denote the set of all permissible tokens associated with a language model. Given a context $x^{t}=(x(1),\ldots,x(t))\in\Omega^{t}$ our target model $s$ produces conditional probabilities ${\mathcal{M}}_{s}(y|x^{t})$ for each $y\in\Omega$. Following prior works and as explicitly stated in Sun et al. (2024b) we assume the following computational model:
  1. 1.
    Standard Inference: Given a context $x^{t}$, an auto-regressive model can output ${\mathcal{M}}_{b}(y|x^{t})$ for each $y\in\Omega$ in $O(1)$ time.
  2. 2.
    Parallelization along time-axis Given a context $x^{t}$, an auto-regressive model can output ${\mathcal{M}}_{b}(y|x^{i})$ for all $i\in\{1,\ldots,t\}$ and each $y\in\Omega$ again in $O(1)$ time.
Since the standard inference is a sequential process, the idea in Leviathan et al. (2023), Chen et al. (2023) is to exploit parallelization along time-axis to speed up inference. This is accomplished using a draft model ${\mathcal{M}}_{s}$ that is capable of generating multiple tokens with a much lower computation than ${\mathcal{M}}_{b}(\cdot)$. The main steps are summarized below. We assume that a context token $x^{t}$ is the input.
  1. 1.
    Draft Construction: The draft model can efficiently sample $L$ tokens in an auto-regressive manner, extending the context $x^{t}$ to $x(1),\ldots,x(t),\tilde{x}(t+1),\ldots,\tilde{x}(t+L)$. In addition, we keep the conditional probabilities ${\mathcal{M}}_{s}(y|x_{1}^{t},\tilde{x}_{t+1}^{t+i})$ for each $0\leq i\leq L$ and each $y\in\Omega$.
  2. 2.
    Conditional Probability Computation: Given the samples $\tilde{x}_{t+1}^{t+i}$ generated by ${\mathcal{M}}_{s}$ we compute the conditional probabilities ${\mathcal{M}}_{b}(y|x_{1}^{t},\tilde{x}_{t+1}^{t+i})$ for each $0\leq i\leq L$ and each $y\in\Omega$ using parallelization along time-axis.
  3. 3.
    Draft Selection We select first $L^{\prime}\leq L$ tokens and set $x(t+i)=\tilde{x}(t+i)$ for $i\leq L^{\prime}$. The selection is based on speculative decoding algorithm below.
Algorithm 1 Speculative Sampling
1:  Input: Distributions $p(\cdot)$ and $q(\cdot)$, Sample $X\sim p(\cdot)$
2:  Compute residual distribution $p^{\mathrm{res}}(x)=\frac{q(x)-\min(p(x),q(x))}{1-\sum_{x^{\prime}}\min(p(x^{\prime}),q(x^{\prime}))}$
3:  Accept = False
4:  With probability $\beta_{p,q}(X)=\min\left(1,\frac{q(X)}{p(X)}\right)$, set Accept = True.
5:  if Accept = True then
6:     $Y=X$
7:  else
8:     $Y\sim p^{res}(\cdot)$
9:  end if
10:  Return $Y$.
For convenience we denote $p(y)\equiv M_{s}(y|x^{t})$ and $q(y)\equiv M_{b}(y|x^{t})$. The selection procedure is as follows: Given a context $x^{t}$ and a candidate draft sequence $\tilde{x}_{t+1}^{t+L}$, the algorithm in the first step sets $p(\tilde{x}_{t+1})={\mathcal{M}}_{s}(\tilde{x}_{t+1}|x_{1}^{t})$ and $q(\tilde{x}_{t+1})={\mathcal{M}}_{b}(\tilde{x}_{t+1}|x_{1}^{t})$ respectively. It accepts $\tilde{x}_{t+1}$ with probability $\beta_{p,q}(\tilde{x}_{t+1})=\min\left(1,\frac{q(\tilde{x}_{t+1})}{p(\tilde{x}_{t+1})}\right)$. If the token is accepted then $x(t+1)=\tilde{x}_{t+1}$ and we proceed with $\tilde{x}_{t+2}$. If the token is rejected then we sample $Y\sim p^{\mathrm{res}}(\cdot)$ and set $x(t+1)=Y$. At this step we initiate a call to sample $L$ fresh tokens from the draft model with $x_{1}^{t+1}$ as input. This process continues until all $L$ tokens have been accepted or a token is rejected. The correctness of the proposed algorithm follows through direct computation of $\Pr(Y=y)$. We also compute the probability of accept:
$\Pr(\mathrm{accept})$ $=\sum_{x\in{\mathcal{X}}}\Pr(\mathrm{acc}|X=x)p(x)$ (13)
$=\sum_{x\in{\mathcal{X}}}\beta_{p,q}(x)p(x)$ (14)
$=\sum_{x\in{\mathcal{X}}}\mathrm{min}(p(x),q(x))=1-d_{TV}(p,q),$ (15)
where $d_{TV}(\cdot)$ is the total variational distance. Note that $\Pr(\mathrm{accept})$ is a key metric that determines the probability that a token from the draft model is accepted. The higher the value of $\Pr(\mathrm{accept})$, the more efficient is the algorithm. In this paper we also study $\Pr(\mathrm{accept})$ in the multi-draft setting.
Appendix B Proof of Theorem 1